OpenStudy (anonymous):

How fast must a ball be rolled along a 70-cm-high table so that when it rolls off the edge it will strike the floor at the same distance (70cm) from the point directly below the table edge.

OpenStudy (anonymous):

solve for t in terms of Vx Vx*t = 1/2gt^2 t= 2Vx /g plug into Vx*t =.7 2Vx^2/g =.7 Vx =sqrt(.7g/2)